Estimate the change in the density of water in ocean at a depth of 400 m below the surface. The density of water at the surface = 1030 kg m-3 and the bulk modulus of water = 2 × 109 N m-2.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. . p0 = m/V0 = m/Vd So, pd/p0 = V0/Vd ….(1) Vol.strain = V0 - Vd/ V0 B = p0gh/( V0 - Vd)/V0 ⇒ 1 - Vd/ V0 = p0gh/B ⇒ vD/v0 = (1- p0gh/B) ….(2) Putting vale of (2) in equation (1), we get Pd/p0 = 1/ p0gh/B ⇒ pd = 1/(1 - p0gh/B )* p0